Journalize the following transactions for Marge's SofasMarge's Sofas.

SOLVE ONLY FOR i. Adjusted for overallocated or underallocated overhead.

Given information

a.Incurred and paid Web site​ expenses $2,000.

b.Incurred manufacturing wages of $15,000​, 75​% of which was direct labor and 25​% of which was indirect labor.

c.Purchased raw materials on​ account $24,000.

d.Used in​ production: direct​ materials,$7,500​ ;indirect​ materials,$5,000.

e.Recorded manufacturing​ overhead: depreciation on​ plant,$18,000​;plant insurance​ (previously paid),$1,500​;plant property​ tax, $3,900​(credit Property Tax​ Payable).

f.Allocated manufacturing overhead to​ jobs, 200​%

of direct labor costs.

g.Completed production on jobs with costs of $40,000.

h.Sold inventory on​ account 22,000​;cost of goods​ sold $ 18,000

i.Adjusted for overallocated or underallocated overhead.

​(Record debits​ first, then credits. Exclude explanations from any journal​ entries.)

a. Incurred and paid Web site​ expenses,$2,000.

  

Date

Accounts

Debit

Credit

a.

Website Expenses

2,000

Cash

2,000

b. Incurred manufacturing wages of $15,000​, 75​% of which was direct labor and 25​% of which was indirect labor.

Date

Accounts

Debit

Credit

b.

Work-in-Process Inventory

11,250

Manufacturing Overhead

3,750

Wages Payable

15,000

c. Purchased raw materials on​ account, $24,000.

Date

Accounts

Debit

Credit

c.

Raw Materials Inventory

24,000

Accounts Payable

24,000

d. Used in​ production: direct​ materials, $7,500​; indirect​ materials, $5,000.

Date

Accounts

Debit

Credit

d.

Work-in-Process Inventory

7,500

Manufacturing Overhead

5,000

Raw Materials Inventory

12,500

e. Recorded manufacturing​ overhead: depreciation on​ plant, $18,000​; plant insurance​ (previously paid), $1,500​; plant property​ tax, $3,900 (credit Property Tax​ Payable). ​(Prepare a compound​ entry.)

Date

Accounts

Debit

Credit

e.

Manufacturing Overhead

23,400

Accumulated Depreciation–Plant

18,000

Prepaid Insurance

1,500

Property Tax Payable

3,900

f. Allocated manufacturing overhead to​ jobs, 200​% of direct labor costs.

Date

Accounts

Debit

Credit

f.

Work-in-Process Inventory

22,500

Manufacturing Overhead

22,500

g. Completed production on jobs with costs of $40,000.

Date

Accounts

Debit

Credit

g.

Finished Goods Inventory

40,000

Work-in-Process Inventory

40,000

h. Sold inventory on​ account, $22,000​; cost of goods​ sold, $18,000.

Begin by recording the revenue from the sales on account. Do not record the expense related to the sale in this journal entry. We will do that in the next step.

Date

Accounts

Debit

Credit

h.

Accounts Receivable

22,000

Sales Revenue

22,000

Now record the expense related to the sale.

Date

Accounts

Debit

Credit

h.

Cost of Goods Sold

18,000

Finished Goods Inventory

18,000

i. Adjusted for overallocated or underallocated overhead.

Date

Accounts

Debit

Credit

i.

Cost of Goods Sold

Manufacturing Overhead
Answer #1

1. Actual Manufacturing OH = 3750 + 5000 + 23400 = $32150

2. Applied Manufacturing OH = 22500

Under Allocated OH = Actual Manufacturing OH - Applied OH

Under Allocated OH = 32150 - 22500 = $9650

Accounts

Accounts

Debit

Credit

Cost of Goods Sold $9650
Manufacturing Overhead $9650